# Zero-Shot Position Debiasing for Large Language Models

### v1

@@ 4.1 Datasets
  • Mutual (Cui et al., 2020) is a multi-turn reasoning dialogue, consisting of 8,860 manually annotated dialogues based on Chinese student English listening comprehension exams. It is challenging since it requires a model to handle various reasoning problems.
@@
4.2 Bias types and Dataset Splitting
@@
  • Lead bias in summarization is that the generated summary is highly correlated to utterances appearing at the beginning of the document (Kedzie et al., 2018).
  • Relative position bias in QA is a phenomenon that a QA model tends to degrade the performance on samples where answers are located in relative positions unseen during training (Shinoda et al., 2022). The relative position is defined as the relative position between utterances containing the last turn answer span and the current turn answer span.
  • Lexical bias is the phenomenon that deep learning models achieve high accuracy by exploiting trigger words or word overlapping (Gururangan et al., 2018; Poliak et al., 2018).
Following previous works (Ko et al., 2020; Shinoda et al., 2022), we split the samples into biased samples and non-biased samples for simulation depending on the bias type in each dataset. In CQG and CQA datasets (CANARD and CoQAR), we select the samples with relative position equaling 0 or 1 as biased samples and the left samples as non-biased samples. In KGC and summarization datasets (Doc2dial, Mutual, CNN/DM and Newsroom), we filter samples where the reference response is highly correlated to the beginning utterance of the given document into biased samples and the left samples into non-biased samples. For NLI datasets (SNLI and QNLI), samples with specific words are filtered into biased samples and other samples are filtered into non-biased samples. The dataset statistics are shown in Table 1.
@@
Dataset Biased
Non-biased
#Train
#Dev
#Test
#Test
CANARD
500
250
3,460
2,440
CoQAR
500
250
3,222
4,873
CNN/DM
500
250
1,421
5,000
Newsroom
500
200
5,000
5,000
Doc2dial
500
250
2,000
5,000
Mutual
500
250
2,000
5,000
SNLI
500
250
2,000
5,000
QNLI
500
250
2,000
5,000
4.3 Evaluation Metrics
Following previous works (Chen et al., 2019; Nallapati et al., 2016; Tuan et al., 2020; Meng et al., 2020), we adopt ROUGE-L Lin (2004) as evaluation metrics for CQA, CQG, summarization and KGC tasks, in which ROUGE-L has been shown to correlate well with human evaluation (Liu and Liu, 2008). We use macro-accuracy for NLI task as NLI is indeed a classification task. We use nlg-eval package [^3] https://github.com/Maluuba/nlg-eval for the implementation of evaluation metrics.
4.4 Baseline Methods
  • Random Position (RT) (Shinoda et al., 2022) randomly perturbs input positions to reduce the model’s dependence on token positions in prediction.
  • FlanT5 (FT) is the FlanT5-large model (Chung et al., 2022) fine-tuned on the dataset for the target task to improve the performance of the target task.
  • Minimax (Korakakis and Vlachos, 2023) leverages an auxiliary model to maximize the learner’s loss by up-weighting ‘hard’ samples, thus reducing the reliance of NLI models on shortcuts in ‘easy’ samples.
4.5 Implementation Details
We use FlanT5-large (Chung et al., 2022) as the base model for all models. The hidden size is 768. We use the Adam optimizer with a default learning rate 1e-4 (Kingma and Ba, 2015) and set gradient clipping with a default maximum gradient norm of 1.0444https://huggingface.co/docs/transformers/v4.33.0/en/main_classes/trainer. We select the best model based on the BLEU@2 or macro-accuracy score on the validation set. We set $\alpha$=0.1 for CQG and CQA and summarization tasks, and $\alpha$=0.2 for NLI and KGC tasks, by default. To facilitate the reproducibility of the results, we will share the code of all methods and datasets soon.
5 Results
@@
Table 2: Overall performance (%) on language understanding tasks. Boldface indicates the best results in terms of the corresponding metrics.
@@
T5 77.5 79.0 90.3 88.3 47.4 39.8 34.0 17.3
RP 61.5 51.6 60.4 20.4
MarCQAp 66.9 52.3 66.3 21.3
Minimax 91.6 87.0 90.9 89.7
FT 91.8 86.6 93.0 89.8 64.6 51.9 67.5 20.8
ZOE 91.8 89.0 93.3 91.3 66.3 53.7 65.7 21.9
Table 3: Overall performance (%) on language creation tasks. Boldface indicates the best results in terms of the corresponding metrics.

### v3

@@ 4.1 Datasets
We conduct experiments on eight widely used benchmark datasets: CANARD (Elgohary et al., 2019), CoQAR (Brabant et al., 2022), CNN/DM (Nallapati et al., 2016), Newsroom (Grusky et al., 2018), Doc2dial (Feng et al., 2020), Mutual (Cui et al., 2020), SNLI (Bowman et al., 2015) and QNLI (Wang et al., 2018), covering five NLP tasks: CQA, CQG, summarization, KGC and NLI. Following previous works (Ko et al., 2020; Shinoda et al., 2022), we split the test dataset into biased dataset and non-biased dataset for simulation depending on the bias type in each dataset. The details of datasets and dataset splitting are provided in §A.1 and §A.2.
4.2 Evaluation Metrics
Following previous works (Chen et al., 2019; Nallapati et al., 2016; Tuan et al., 2020; Meng et al., 2020), we adopt ROUGE-L Lin (2004) as evaluation metrics for CQA, CQG, summarization and KGC tasks, in which ROUGE-L has been shown to correlate well with human evaluation (Liu and Liu, 2008). We use macro-accuracy for the classification task, NLI. We use nlg-eval package [^3] https://github.com/Maluuba/nlg-eval for the implementation of evaluation metrics.
4.3 Baseline Methods
  • BASE is the pre-trained LLM with unsupervised instruction-following fine-tuning.
  • Random Position (RP) (Shinoda et al., 2022) randomly perturbs input positions to reduce the model’s dependence on token positions in prediction.
  • Fine-tune (FT) is the LLM fine-tuned on the dataset for the target task to improve the performance of the target task.
  • Minimax (Korakakis and Vlachos, 2023) is an NLI model which leverages an auxiliary model to maximize the loss of the NLI model by up-weighting ‘hard’ samples, thus reducing its reliance of shortcuts in ‘easy’ samples.
4.4 Implementation Details
We use FlanT5-large (Chung et al., 2022) as the base LLM for all models. The hidden size is 768. We use the Adam optimizer with a default learning rate $1e^{-4}$ (Kingma and Ba, 2015) and set gradient clipping with a default maximum gradient norm of 1.0444https://huggingface.co/docs/transformers/v4.33.0/en/main_classes/trainer. We select the best model based on the BLEU@2 or macro-accuracy score on the validation set. We use $\alpha$=0.2 for CQA on CoQAR, NLI and KGC tasks and $\alpha$=0.1 for other tasks, by default. We set the pre-defined thresholds for incoherent identification and unreliable identification from 0.1, 0.15 and 0.2 and select the one that maintains approximately 20% unsupervised responses. We run all experiments with NVIDIA RTX3090 24 GB GPU cards.
5 Results
Table 1: Overall performance (%) on language understanding tasks. Boldface indicates the best results in terms of the corresponding dataset.
Method NLI (%) CQA (%)
SNLI QNLI CoQAR CANARD
BASE 77.5 79.0 90.3 88.3 47.4 39.8 34.0 17.3
RP 61.5 51.6 60.4 20.4
MarCQAp 66.9 52.3 66.3 21.3
Minimax 91.6 87.0 90.9 89.7
FT 92.0 88.0 94.3 89.8 64.6 51.9 67.5 20.8
SOD 92.0 88.4 94.3 90.9 66.3 53.7 65.7 21.9
Table 2: Overall performance (%) on language creation tasks. Boldface indicates the best results in terms of the corresponding dataset.
Method CQG (%) KGC (%)
CoQAR CANARD Doc2dial Mutual
BASE 16.0 17.0 17.9 17.1 23.9 13.7 25.4 21.3
RP 23.2 18.3 24.6 21.2 35.3 32.2 89.6 46.1
MarCQAp 19.7 17.5 26.0 21.8
SG-CQG 14.9 15.5 19.7 17.8
FocusL 39.8 35.9 83.0 51.7
FT 26.7 17.2 26.0 21.6 38.7 36.0 93.9 38.9
SOD 26.7 18.8 25.9 22.4 40.6 38.3 94.0 53.0
Table 3: Overall performance (%) on language compression tasks. Boldface indicates the best results in terms of the corresponding dataset.
Method Summarization (%)
CNN/DM Newsroom
Biased Non-Biased Biased Non-Biased
BASE 22.3 11.6 35.1 20.0
RP 23.9 15.1 47.5 22.0
GenX 17.6 13.7 29.2 19.0
FT 26.9 16.8 51.1 19.8
SOD 27.1 17.3 50.9 21.3
The overall performances of all methods on language understanding tasks, language creation tasks and language compression tasks are listed in Table 1-3. We have three main observations from the results.
First, LLMs are susceptible to the bias in the dataset after fine-tuning. As we can see in Table 1, FT achieves 34.7% improvement on the biased dataset of CoQAR, but 8.6% improvement on the non-biased dataset. This is because LLMs can easily overfit the shortcut of the training dataset in fine-tuning, just like existing neural networks (Ko et al., 2020).
Second, SOD can mitigate position bias significantly on almost all datasets for three types of tasks. As shown in Table 1-3, SOD improves the performance on the non-biased dataset by 1% to 2% on almost all tasks, compared to FT and all baselines. The reason is that SOD can leverage unsupervised responses with low position bias for optimization in multi-objective optimization module.
Third, SOD only sacrifices a small performance on the biased dataset when mitigating position bias. As shown in Table 3, RP achieves comparable performance to SOD on the non-biased dataset of Newsroom. However, the ROUGE-L of RP drops 3.6% compared to FT on the biased dataset, while that of SOD only drops 0.2%. This is because the perturbation in RP impairs the overall data quality for fine-tuning, while unsupervised responses in SOD are aligned to improve the quality in §3.2.
Note that some baselines achieve poor performance, sometimes lower than BASE. First, in Table 3, GenX performs worse even than BASE on the biased dataset of Newsroom. The reason is that the summarization datasets are abstractive and suitable for generative models, e.g., T5, while GenX is an extractive baseline. Second, in Table 2, the performance of SG-CQG is worse than that of FT on the biased dataset of CoQAR. This is because SG-CQG is a three-stage question generation model focusing on improving question diversity. The selected answer span for question generation is randomly chosen from massive generated candidates.
Besides, to further verify the effectiveness of SOD, we have done t-test significant tests. We found that SOD can achieve significant improvement over MarCQAp on CoQAR (CQA and CQG) and CANARD (CQG), over FocusL on Doc2dial and Mutual, over FT on CNN/DM. The above results indicate that SOD can mitigate position bias in most cases.
6 Analysis
In this section, we analyze the effect of the quality of unsupervised responses in §6.1 and objective weighting in the multi-objective optimization module in §6.2. The overall results of all tasks are presented in §C. Besides, we also conduct a case study in §B.1 and provide cases for all datasets in §C.4.
6.1 Analysis of OAM
To analyze the effect of the OAM module, we conduct analyses with unsupervised responses obtained from various sources with different qualities in Table 4. SOD w/o OAM, SOD w/ T5-base and SOD w/ T5-xlarge denote SOD using unsupervised responses without alignment, responses from FlanT5-base and FlanT5-xlarge, respectively. We have two observations.
First, low-quality responses without OAM leads to worse performance. As we can see, SOD outperforms SOD w/o OAM on non-biased datasets by leveraging OAM for enhancing the response quality. Poor-quality responses will undermine the model comprehension of the task, thus leading to worse performance.
Table 4: SOD performance (%) of CQG task using various unsupervised responses. ‘N-Biased’ denotes performance on non-biased datasets. Boldface indicates the best results in terms of the corresponding dataset.
Method CoQAR CANARD
Biased N-Biased Biased N-Biased
SOD 26.7 18.8 25.9 22.4
- w/o OAM 25.8 18.0 26.1 21.5
- w/ T5-base 26.3 18.3 26.1 22.0
- w/ T5-xlarge 26.6 17.9 25.7 22.0
FT 26.7 17.2 26.0 21.6
Second, OAM is robust to various sources of unsupervised responses. SOD w/ T5-base and SOD w/ T5-xlarge perform worse than SOD on CANARD, but still outperform FT. We infer that responses from other LLMs use different knowledge/parameters for generation, which mismatch with that of T5. The difference amplifies the divergence between the task objective and the debiasing objective in §3.2, thus leading to worse performance. Even that, OAM can reduce this divergence to better achieve both objectives.
6.2 Analysis of Objective Weighting
[CoQAR (CQA)]Refer to caption [CANARD (CQG)]Refer to caption
[Mutual]Refer to caption [QNLI]Refer to caption
Figure 3: Performance (%) of four tasks over each $\alpha$. The x-axis denotes the value of $\alpha$ and the y-axis denotes the ROUGE-L score on non-biased datasets.
To analyze the effect of weighting on the debiasing objective, we present the performance of SOD using different $\alpha$ in Fig. 3. We have two observations from the results.
First, the performance of SOD drops with the increase of the weight of unsupervised responses in multi-objective optimization. In CQA on CoQAR, the ROUGE-L score of SOD drops from 53.6% to 52.7% when increasing $\alpha$ from 0.1 to 0.5. This is because the model performance depends on not only the degree of bias but also the data quality. Increasing $\alpha$ will reduce the position bias of all responses, yet it will hurt the quality concurrently.
Second, our proposed SOD always outperforms FT under various $\alpha$. As shown in Fig. 3, SOD performances of CQA on CoQAR all exceed 52.5% using various $\alpha$, while FT only achieves 52.0%. This demonstrates the effectiveness and robustness of SOD in mitigating position bias.
6.3 Analysis on Training Samples
[CoQAR (CQA)]Refer to caption [CANARD (CQG)]Refer to caption
[Mutual]Refer to caption [QNLI]Refer to caption
Figure 4: Performance (%) of four tasks over different numbers of training samples. The x-axis denotes the number of training samples and the y-axis denotes the ROUGE-L score on non-biased datasets.
We also analyze the effect of the number of training samples to verify the effectiveness of SOD under various low-resource settings. We plot the results in Fig. 4 and have two observations.
First, nearly all methods perform better when increasing training samples. As we can see in Fig. 4, the ROUGE-L score of FT increases from 19.7% to 22.5% on CANARD when the number of training samples increases from 50 to 1,000. Increasing training samples can improve the generalization ability of LLMs, thus leading to better performance.
Second, our proposed SOD outperforms FT under various low-resource settings. As shown in Fig. 4, the ROUGE-L scores of SOD depicted by the orange bars are consistently higher than those of FT in blue. This is because there are always around 40% aligned unsupervised responses for fine-tuning regardless of the variation in the number of training samples, which are enough for effective debiasing.
7 Related Work
7.1 Position Bias and Debiasing
Position bias has been explored mainly in three extractive tasks, i.e., NLI, summarization and CQA. In, NLI, Gururangan et al. (2018); Poliak et al. (2018) show that the class labels are highly correlated to certain words in the hypothesis. McCoy et al. (2019) report that models always rely on word overlap between hypothesis and premise for prediction. Karimi Mahabadi et al. (2020); Du et al. (2021) train a robust NLI model in an ensemble manner by PoE. They train a hypothesis model to learn the lexical bias, which guides the NLI model to focus on other patterns in the dataset that generalize better. Ghaddar et al. (2021) re-weight the importance of easy and hard samples to prevent the model from fitting the bias. As the weights are derived from the model itself, this method does not require any external annotations.
Similar findings are reported in CQA tasks. Weissenborn et al. (2017); Sugawara et al. (2018) demonstrate that only using partial inputs is sufficient to correctly extract the answer span for the question in most cases. Ko et al. (2020) address that absolute position of answer spans can work as a spurious clue for prediction. Shinoda et al. (2022) report that relative position of answer spans is another clue for position bias. To mitigate position bias, Ko et al. (2020); Shinoda et al. (2022) build bias ensemble models by PoE similar to Karimi Mahabadi et al. (2020). They design biased models with position-only features to guide CQA models to rely more on semantic features for answering.
Kedzie et al. (2018); Grenander et al. (2019) find that 58% of selected summary utterances come directly from the lead utterances, and models trained on these articles perform considerably worse when utterances in the article are randomly shuffled. To mitigate lead bias, a simple but effective method is to randomly shuffle the document for training (Grenander et al., 2019). Then, Xing et al. (2021) uses adversarial training for debiasing. They design a position prediction module and optimize the reverse loss for position prediction, forcing the encoder to leverage non-position features.
However, existing works on mitigating position bias always focus on extractive tasks. They can hardly be transferred to generative tasks as the label space in generative tasks is too large for bias estimation or adversarial training. In this paper, we focus on position debiasing for generative LLMs.
7.2 Debiasing for LLMs
In pre-processing, Zmigrod et al. (2019) adopt a counterfactual data augmentation (CDA) algorithm to mitigate social bias by swapping bias attribute words (e.g., he/she) in training dataset. Choi et al. (2022) modify CDA by masking the terms casual to label to force the model to learn label-invariant features.
In in-processing, Guo et al. (2022); Li et al. (2023); Yang et al. (2023) propose a two-stage adversarial method for debiasing. They first train a continuous prompt to enlarge the bias of utterance pairs and then force the LLMs to minimize the difference of utterance pairs using the prompt.
In post-processing, Schick et al. (2021) propose a decoding algorithm that reduces the probability of a model producing biased text. They use a textual description of the undesired behaviors for prompting. Meade et al. (2023) propose an ICL strategy which leverages non-biased demonstrations to guide the generation for safety.
However, existing methods for mitigating social bias either require the external bias knowledge for data augmentation and training adversarial prompts or require a non-biased dataset for building demonstrations, which are lacking for position debiasing and unpractical in application. Differently, we propose a self-supervised framework for LLMs on mitigating position bias, without relying on any external bias knowledge or non-biased samples, which is general, simple but effective.
8 Conclusion
In this paper, we have proposed a self-supervised debiasing framework SOD for LLMs. It adopts a multi-objective optimization module to mitigate position bias for LLMs, where unsupervised responses for the debiasing objective are of low position bias. These responses are pruned by a proposed OAM module for aligning the task and debiasing objectives. Extensive experiments on five tasks and eight benchmark datasets show that SOD outperforms existing baselines on non-biased samples while retaining performance or sacrificing little performance on biased samples. It demonstrates that leveraging unsupervised responses is a practicable solution to mitigate position bias for generative LLMs.
Limitations
This work has the following limitations. First, SOD needs a pre-trained LLM to generate unsupervised responses with low bias, where these responses are still biased. Second, the final performance of SOD depends on the quality of unsupervised responses, which are still noisy after being aligned by OAM. In future work, we plan to address these issues by investigating non-biased models from other domains and model-based strategies to align unsupervised responses.
Ethical Considerations
We realize that there are risks in developing generative LLMs, so it is necessary to pay attention to the ethical issues of LLMs. We use publicly available pre-trained LLMs, i.e., FlanT5-base, FlanT5-large, FlanT5-xlarge, and publicly available datasets in the academic community, i.e., CANARD, CoQAR, CNN/DM, Newsroom, Doc2dial, Mutual, SNLI, QNLI, to conduct experiments. All models and datasets are carefully processed by their publishers to ensure that there are no ethical problems.
Acknowledgments
We thank the reviewers for their valuable feedback. This work was supported by the National Key R&D Program of China with grant No.2022YFC3303004, the Natural Science Foundation of China (62372275, 62272274, 62202271, T2293773, 62102234, 62072279), the Natural Science Foundation of Shandong Province (ZR2021QF129).
Appendix A Experimental Details
A.1 Datasets
@@
  • Mutual (Cui et al., 2020) is a multi-turn reasoning dialogue dataset, consisting of 8,860 manually annotated dialogues based on Chinese student English listening comprehension exams. It is challenging since it requires a model to handle various reasoning problems.
@@
A.2 Bias Types and Dataset Splitting
@@
  • Lead bias in summarization is a phenomenon that the generated summary is highly correlated to utterances appearing at the beginning of the document (Kedzie et al., 2018).
  • Relative position bias in QA is a phenomenon that a QA model tends to degrade the performance on samples where answers are located in relative positions unseen during training (Shinoda et al., 2022). The relative position is defined as the relative position of grounded utterances between the last turn answer and the current turn answer.
  • Lexical bias is the phenomenon that deep learning models achieve high accuracy by exploiting trigger words or word overlapping (Gururangan et al., 2018; Poliak et al., 2018). Note that lexical bias is a type of position bias in generative models. For example, trigger words in hypothesis in generative models are regarded as trigger words behind ‘Hypothesis: ’, which is a positional clue for prediction.
Following previous works (Ko et al., 2020; Shinoda et al., 2022), we split the test dataset into biased dataset and non-biased dataset for simulation depending on the bias type in each dataset. In CQG and CQA datasets (CANARD and CoQAR), we select the samples with relative position equaling 0 or 1 into the biased dataset and the left samples into the non-biased dataset. In KGC and summarization datasets (Doc2dial, Mutual, CNN/DM and Newsroom), we filter samples where the reference response is highly correlated to the beginning utterance of the given document into the biased dataset and the left samples into the non-biased dataset. For NLI datasets (SNLI and QNLI), samples with specific words are filtered into the biased dataset and other samples are filtered into the non-biased dataset. The dataset statistics are shown in Table 5.
@@
Dataset Biased Non-biased
#Train #Dev #Test #Test
CANARD 500 250 3,460 2,440
CoQAR 500 250 3,222 4,873
CNN/DM 500 250 1,421 5,000
Newsroom 500 200 5,000 5,000
Doc2dial 500 250 2,000 5,000
Mutual 500 250 2,000 5,000
SNLI 500 250 2,000 5,000
QNLI 500 250 2,000 5,000
Appendix B Analysis
B.1 Case Study
Table 6: An example of generated answers on CANARD. U1, U4, and U5 are the 1st, 4th, 5th utterances in the document, T1 and T2 are the first and second turn utterances of the dialogue. ‘Target’ is the target answer of this example, ‘FT’ and ‘SOD’ are the generated answers of FT and SOD. ‘Position’ is the position of the grounded utterance of the answer in the document and ‘Relative position’ is the distance of grounded utterances between the current answer and the last turn answer.
ID Document
U1 Gaston had worked with players at an individual level as a hitting instructor …
U4 In the six games the Blue Jays played in those places during World Series play
U5 … and Gaston was the first ever African-American manager to win a World Series.
ID Context Position
T1 Question: What happens in the series?
U4
T2 Question: What else did Cito do?
Model Answer Relative position
Target Gaston had worked with players at an individual level -3 (U1)
FT Gaston was the first ever African-American manager to win a World Series.  1 (U5)
SOD Gaston had worked with players at an individual level -3 (U1)
Table 7: An example of unsupervised questions on CANARD. U1–U6 are the first six utterances in the document and T1 is the first turn utterance of the dialogue. ‘Target’ is the target question of this example, ‘BASE-1’ and ‘BASE-2’ are the questions generated by pre-trained LLM. ‘Position’ is the position of the grounded utterance of the question in the document and ‘Relative position’ is the distance of grounded utterances between the current question and the last turn question.
ID Document
U1 Gautam Gambhir, born 14 October 1981, is an Indian cricketer, …
U2 Gambhir was picked up by the Delhi Daredevils franchise in the first player auction of the Indian Premier League for a price of US\$725,000 a year.
U3 He became the second highest run-scorer of the inaugural season with …
U4 He was promoted to the post of Captain of the Delhi Daredevils for IPL Season 2010.
U5 At the end of the tournament he became the only player from Delhi Daredevils to score more than 1000 runs in the IPL.
U6 In the 2011 IPL player auction, …
ID Context Position
T1 Question: Who was Gautam Gambhir in Indian premier league?
U4
Source Question Relative position
Target What did Gautam Gambhir do as captain?  1 (U5)
BASE-1 Where was Gautam Gambhir born? -3 (U1)
BASE-2 Was Gautam Gambhir in Indian premier league? -2 (U2)
To investigate the reason for the effectiveness of SOD, we present an example of generated responses in CQA in Table 6 and an example of unsupervised responses in CQG in Table 7.
In CQA example, SOD can generate an answer with lower bias than FT. As shown in Table 6, FT generates an answer from U5, adjacent to U4, the utterance containing the last turn answer. In contrast, SOD generates an answer from U1, which is far from U4. After fine-tuning with unsupervised responses from various positions, SOD cannot easily fall into the trap of finding answers from neighboring utterances of the one containing the last turn answer.
In CQG example in Table 7, SOD uses unsupervised responses which have lower bias than the target response. As we can see, target question is based on the utterance with relative position 1, while unsupervised questions generated from BASE are based on utterances varying in the document. Fine-tuning with unsupervised responses generated from pre-trained LLMs encourages the model to generate questions based on utterances in various document positions, thus mitigating position bias.
Appendix C Overall Analysis
@@
Table 8: SOD Performance (%) using various unsupervised responses on language understanding tasks. Boldface indicates the best results in terms of the corresponding dataset.
@@
SOD 92.0 88.4 94.3 90.9 66.3 53.7 65.7 21.9
- w/o OAM 91.0 87.4 92.7 88.7 66.3 53.7 65.7 21.9
- w/ T5-base 90.5 87.6 94.4 90.6 64.3 52.3 61.6 19.9
- w/ T5-xlarge 91.2 87.7 93.2 91.3 65.1 53.6 67.1 22.4
FT 92.0 88.0 94.3 89.8 64.6 51.9 67.5 20.8
Table 9: SOD Performance (%) using various unsupervised responses on language creation tasks. Boldface indicates the best results in terms of the corresponding dataset.